# Approximating Two-Layer ReLU Networks for Hidden State Analysis in Differential Privacy

### v2

@@ Appendix D Utility Bound within the Random Data Model
$M=\mathbb{E}_{g\sim\mathcal{N}(0,I_{d})}[\mathrm{diag}[\mathds{1}(Xg\geq 0)]XX^{T}\mathrm{diag}[\mathds{1}(Xg\geq 0)]].$
@@
There results together tell that taking $d$ and $n$ large enough (such that $n\geq d$), we have that with $P={O}(\frac{n\log\frac{n}{\gamma}}{d})$ hyperplane arrangements we get the zero global optimum with high probability, i.e., there exists $u\in\mathbb{R}^{d\cdot P}$ such that Eq. (D.1) holds with probability at least $1-\gamma-\frac{1}{(2n)^{8}}$.
In case $d$ and $n$ are large enough and we choose $P={O}\left(\frac{n\log\frac{n}{\gamma}}{d}\right)$ hyperplane arrangements, we have that $p=d\cdot P={O}(n\log\frac{n}{\gamma})$ and we directly get the following corollary.
We can directly apply the following classical result from ERM for the DP-GD (Alg. 1) to the stochastic problem (3.4) or (3.6).
Theorem D.3 (Bassily et al., 2014; Talwar et al., 2014).
Assuming the gradients stay bounded by a constant $L$, this result gives utility bounds for the stochastic problems (3.4) or (3.6) with $p=d\times P$.
@@
For any $\gamma>0$, there exists $d_{1}$ such that for all $d\geq d_{1}$, $n\geq d$, with probability at least $1-\gamma-\frac{1}{(2n)^{8}}$,
$\mathcal{L}(\theta^{\textrm{priv}},X)\leq\widetilde{O}\left(\frac{1}{\sqrt{n}\varepsilon}\right),$
where $\widetilde{O}$ omits logarithmic factors.
@@
By Lemma D.2, with with probability at least $1-\gamma-\frac{1}{(2n)^{8}}$, for $p=d\cdot P={O}(n\log\frac{n}{\gamma})$, we have that $\mathcal{L}(\theta^{*},X)=0$. Substituting this $p$ to the claim of Theorem D.3, the claim follows. ∎
@@ Appendix E Illustrations of the Stochastic Approximation
E.1 Illustration with SDG Applied to MNIST
@@
Figure 4 illustrates the approximability of the stochastic approximation for the dual problem in the non-private case, when the number of random hyperplanes $P$ is varied, for the MNIST classification problem described in Section 6. We apply SGD with batch size 1000 to both the stochastic dual problem and to a fully connected ReLU network with hidden-layer width 200, and for each model optimize the learning rate using the grid $\{10^{-i/2}\}$, $i\in\mathbb{Z}$. This comparison shows that the approximabilty of the stochastic dual problem increases with increasing $P$.
@@
Figure 4: Test accuracies vs. number of epochs, when all models are trained using SGD with batch size 1000. The number of random hyperplanes $P$ is varied for the stochastic dual problem. The ReLU network is a one hidden-layer fully connected ReLU network with hidden-layer width 200. Cross-entropy loss is used for all models.
E.2 Illustration with DP-SDG Applied to MNIST
@@
Figure 5 illustrates the approximability of the stochastic approximation for the dual problem in the private case, when the number of random hyperplanes $P$ is varied, for the MNIST classification problem described in Section 6. We apply DP-SGD with batch size 1000 to both the stochastic dual problem and to a fully connected ReLU network with hidden-layer width 200, and for each model optimize the learning rate using the grid $\{10^{-i/2}\}$, $i\in\mathbb{Z}$. Based on these comparisons, we conclude that $P=128$ is not far from optimum, as increasing the dimension starts means that the adverse effect of the DP noise becomes larger.
@@
Figure 5: Test accuracies vs. number of epochs, when all models are trained using DP-SGD with batch size 1000, for two different noise levels $\sigma$. The number of random hyperplanes $P$ is varied for the stochastic dual problem. The ReLU network is a one hidden-layer fully connected ReLU network with hidden-layer width 200.
Appendix F Comparison of PLD and RDP Accounting for Subsampling Without Replacement
Instead of using the numerical approach described in Section 2 of the main text, we could alternatively compute the $(\varepsilon,\delta)$-DP guarantees for DP-SGD with subsampling without replacement using the RDP bounds given by Wang et al., (2019). Fig. 6 illustrates the differences when $\sigma=5.0$ and ratio of the batch size and total dataset size $m/n$ equals 0.01. The RDP parameters are converted to $(\varepsilon,\delta)$-bounds using the conversion formula of Lemma F.1.
Lemma F.1 (Canonne et al., 2020).
Suppose the mechanism $\mathcal{M}$ is $\big{(}\alpha,\epsilon^{\prime}\big{)}$-RDP. Then $\mathcal{M}$ is also $(\epsilon,\delta(\epsilon))$-DP for arbitrary $\epsilon\geq 0$ with
$\delta(\epsilon)=\frac{\exp\big{(}(\alpha-1)(\epsilon^{\prime}-\epsilon)\big{)}}{\alpha}\left(1-\frac{1}{\alpha}\right)^{\alpha-1}.$ (F.1)
Figure 6: $(\varepsilon,\delta)$-DP guarantees for DP-SGD with subsampling without replacement, computed using the RDP bound of Wang et al., (thm. 9, 2019) and the numerical dominating pair computed using the privacy profile bound of Zhu et al., (2022) and the numerical algorithm by Doroshenko et al., (2022). The parameter $\sigma=5.0$ and ratio of the batch size and total dataset size $m/n$ equals 0.01.
Appendix G Reformulation Towards a Dual Form for the Practical Model
An interesting question is whether we can interpret the loss function (3.6) suitable for DP analysis as a stochastic approximation of a dual form of some ReLU minimization problem, similarly as the stochastic problem (3.4) approximates the convex problem  (3.3). We have the following result which is analogous to the reformulation behind the non-strongly convex dual form (3.3). We leave as a future work to find out whether we can state the loss function (3.6) as an approximation of some dual form.
Theorem G.1.
For a data-matrix $X\in\mathbb{R}^{n\times d}$, label vector $y\in\mathbb{R}^{n}$ and a regularization parameter $\lambda>0$, consider the ReLU minimization problem
$\min_{\{u_{j},\alpha_{j}\}_{j=1}^{m}}\frac{1}{2}\left\lVert\sum\nolimits_{j=1}^{m}\phi(Xu_{j})\alpha_{j}-y\right\rVert_{2}^{2}+\frac{\lambda}{2}\left(\sum\nolimits_{j=1}^{m}\left\lVert u_{j}\right\rVert_{2}^{4}+\alpha_{j}^{4}\right).$ (G.1)
Then, the problem (G.1) and the problem
$\min_{\{u_{j},\alpha_{j}\leq 1\}_{j=1}^{m},\left\lVert u_{j}\right\rVert_{2}\forall j\in[m]}\frac{1}{2}\left\lVert\sum\nolimits_{j=1}^{m}\phi(Xu_{j})\alpha_{j}-y\right\rVert_{2}^{2}+\lambda\left(\sum\nolimits_{j=1}^{m}\alpha_{j}^{2}\right)$
have equal minima.
Proof.
From Young’s inequality $\left\lVert x\right\rVert_{2}^{2}+\left\lVert y\right\rVert_{2}^{2}\geq 2\langle a,b\rangle$ it follows that
$\frac{\lambda}{2}\left(\sum\nolimits_{j=1}^{m}\left\lVert u_{j}\right\rVert_{2}^{4}+\alpha_{j}^{4}\right)\geq\lambda\sum\nolimits_{j=1}^{m}\left\lVert u_{j}\right\rVert_{2}^{2}\cdot\alpha_{j}^{2}.$
We see that the problem (G.1) is scaling invariant, i.e., for any solution $\{u_{j}^{*},\alpha_{j}^{*}\}_{j=1}^{m}$ and for any $\gamma_{i}>0$, $i\in[m]$, also $\{u_{j}^{*}\cdot\gamma_{i},\alpha_{j}^{*}/\gamma_{i}\}_{j=1}^{m}$ gives a solution. Choosing for every $i\in[m]$, $\gamma_{i}=\sqrt{\frac{\alpha_{i}}{\left\lVert u_{j}\right\rVert_{2}}}$ gives an equality in Young’s inequality. Since this scaling does not affect the solution, we must have for the global minimizer $\{u_{j}^{*},\alpha_{j}^{*}\}_{j=1}^{m}$ of the ReLU minimization problem (G.1) that
$\frac{1}{2}\left\lVert\sum\nolimits_{j=1}^{m}\phi(Xu_{j}^{*})\alpha_{j}^{*}-y\right\rVert_{2}^{2}+\frac{\lambda}{2}\left(\sum\nolimits_{j=1}^{m}\left\lVert u_{j}^{*}\right\rVert_{2}^{4}+(\alpha_{j}^{*})^{4}\right)$ (G.2)
$=$ $\frac{1}{2}\left\lVert\sum\nolimits_{j=1}^{m}\phi(Xu_{j}^{*})\alpha_{j}^{*}-y\right\rVert_{2}^{2}+\lambda\left(\sum\nolimits_{j=1}^{m}\left\lVert u_{j}^{*}\right\rVert_{2}^{2}(\alpha_{j}^{*})^{2}\right).$
Again, due to the scaling invariance, we see that the minimizing the right-hand-side of (G.2) w.r.t. $\{u_{j},\alpha_{j}\}_{j=1}^{m}$ is equivalent to the problem (G.1).
Appendix H Hyperparameter Grids Used for the Experiments
@@
The hyperparameter grids for the number of random hyperplanes $P$ for the convex model and the hidden width $W$ for the ReLU network are chosen based on the GPU memory of the available machines. For MNIST and FashionMNIST we tune the number of random hyperplanes using the grid
$\{64,128,256\}$
and for CIFAR10 using the grid
$\{16,32,64\}.$
For MNIST and FashionMNIST we tune the hidden width $W$ of the ReLU network using the grid
$\{200,500,800\}$
and for CIFAR10 using the grid
$\{200,400,600\}.$
@@
$\{10^{-3.0},10^{-2.5},10^{-2.0},10^{-1.5},10^{-1.0},10^{-0.5}\}.$
Appendix I Additional Experimental Results on FashionMNIST
Figure 7 shows the accuracies of the best models along the the training iteration of 400 epochs for the FashionMNISTS experiment.
@@ Appendix I Additional Experimental Results on FashionMNIST
Figure 7: FashionMNIST Comparisons: Test accuracies vs. the spent privacy budget $\varepsilon$, when $\delta=10^{-5}$ and each model is trained for 400 epochs. The model is a one hidden-layer fully connected ReLU network and the batch size equals 1000 for all methods considered.

### v3

@@ D.2 Utility Analysis
$\Sigma=\mathbb{E}_{g\sim\mathcal{N}(0,I_{d})}[\mathrm{diag}[\mathds{1}(Xg\geq 0)]XX^{T}\mathrm{diag}[\mathds{1}(Xg\geq 0)]].$
Let $\lambda_{\max}$ and $\lambda_{\min}$ denote the largest and smallest eigenvalues of a matrix, respectively. Kim and Pilanci, (2024) provide the following lower bound on the number of random hyperplanes required, in terms of the condition number $\kappa$, for the problem (D.1) to attain a zero minimum.
@@
Combining Lemmas D.2 and D.2 gives us a high-probability lower bound for the number of hyperplanes in random data model, for the problem (D.1) to attain a zero minimum. Assuming the gradients stay bounded by a constant $L$, by applying the DP-ERM result of Thm.  D.1 gives us the main result (Theorem 4.1 in the main paper):
@@
Consider applying the private gradient descent (Alg. 1) to the strongly convex loss (3.6) and assume that the gradients are bounded by a constant $L>0$. Let the ratio $c=\frac{n}{d}\geq 1$ be fixed. For any $\gamma>0$, there exists $d_{1}$ such that for all $d\geq d_{1}$, with probability at least $1-\gamma-\frac{1}{(2n)^{8}}$ (with $\widetilde{O}$ omitting the logarithmic factors)
$\mathcal{L}(\theta^{\textrm{priv}},D)\leq\widetilde{O}\left(\frac{1}{\sqrt{n}\varepsilon}\right)$
@@
From Lemmas D.2 and D.3 it follows that by taking $d$ and $n$ large enough (such that $n\geq d$), we have that with $P={O}(\frac{n\log\frac{n}{\gamma}}{d})$ hyperplane arrangements there exists $u\in\mathbb{R}^{d\cdot P}$ such that Eq. (D.2) holds with probability at least $1-\gamma-\frac{1}{(2n)^{8}}$. Applying Thm. D.1 with the ambient dimension $p=d\cdot P={O}(n\log\frac{n}{\gamma})$ then shows the claim. ∎
@@ Appendix E Illustrations of the Stochastic Approximation
E.1 Illustration with SGD Applied to MNIST
@@
Figure 4 illustrates the approximability of the stochastic approximation for the dual problem in the non-private case, when the number of random hyperplanes $P$ is varied, for the MNIST classification problem described in Section 5. We apply SGD with batch size 1000 to both the stochastic dual problem and a fully connected ReLU network with hidden-layer width 200. For each model, we optimize the learning rate using the grid $\{10^{-i/2}\}$, $i\in\mathbb{Z}$. This comparison shows that the approximability of the stochastic dual problem increases with increasing $P$.
@@
Figure 4: Test accuracies vs. number of epochs, when all models are trained using SGD with batch size 1000. The number of random hyperplanes $P$ varies for the stochastic dual problem. The ReLU network is a 2-layer fully connected ReLU network with a hidden-layer width of 200. Cross-entropy loss is used for all models.
E.2 Illustration with DP-SGD Applied to MNIST
@@
Figure 5 illustrates the approximability of the stochastic approximation for the dual problem in the private case, when the number of random hyperplanes $P$ is varied, for the MNIST classification problem described in Section 5. We apply DP-SGD with batch size 1000 to both the stochastic dual problem and to a fully connected ReLU network with hidden-layer width 200, and for each model, optimize the learning rate using the grid $\{10^{-i/2}\}$, $i\in\mathbb{Z}$. Based on these comparisons, we conclude that $P=128$ is not far from optimum, as increasing the dimension means that the adverse effect of the DP noise becomes larger.
@@
Figure 5: Test accuracies vs. number of epochs, when all models are trained using DP-SGD with batch size 1000, for two different noise levels $\sigma$. The number of random hyperplanes $P$ varies for the stochastic dual problem. The ReLU network is a 2-layer fully connected ReLU network with hidden-layer width 200.
Appendix F DP Hyperparameter Tuning for DP-SGD and NoisyCGD
When comparing DP optimization methods empirically, it is also important to consider the effect of the hyperparameter tuning on the privacy costs. Most relevant to our work are the private selection methods given by Liu and Talwar, (2019) and Papernot and Steinke, (2022) that are applicable to private hyperparameter tuning and the privacy profile-based analysis of those methods by Koskela et al., (2024). These results apply to tuning algorithms that return the best model of $K$ randomly chosen alternatives, where $K$ is also a random variable. We consider the case where $K$ is Poisson distributed. However, other alternatives exist that allow adjusting the balance between compute cost of training, privacy and accuracy (Papernot and Steinke, , 2022; Koskela et al., , 2024). The DP bounds for Poisson distributed $K$ can be described as follows. Let $Q(y)$ be the density function of the quality score of the base mechanism ($y$ denoting the score) and $A(y)$ be the density function of the tuning algorithm that outputs the best model of the $K$ alternatives. Let $A$ and $A^{\prime}$ denote the output distributions of the tuning algorithm evaluated on neighboring datasets $D$ and $D^{\prime}$, respectively. Then, the hockey-stick divergence between $A$ and $A^{\prime}$ can be bounded using the following result.
Theorem F.1 (Koskela et al., 2024).
$H_{{\rm e}\hskip 1.0pt^{\varepsilon}}(A||A^{\prime})\leq m\cdot\delta(\widehat{\varepsilon}),$ (F.1)
Theorem F.1 says that the hyperparameter tuning algorithm is $\big{(}\widehat{\varepsilon},m\cdot\delta(\widehat{\varepsilon})\big{)}$-DP in case the base mechanism is $(\varepsilon_{1},\delta(\varepsilon_{1})$-DP. If we can evaluate the privacy profile for different values of $\varepsilon$, we can also optimize the upper bound (F.1). When comparing DP-SGD and NoisyCGD, we use the fact that DP-SGD privacy profiles are approximately those of the Gaussian mechanism for large compositions (Sommer et al., , 2019) as follows. Suppose DP-SGD is $(\varepsilon^{*},\delta^{*})$-DP for some values of the batch size $b$ noise scale $\sigma$ and number of iterations $T$. Then, we fix the value of the constant $c=\eta\cdot\lambda$ for NoisyCGD such that it is also $(\varepsilon^{*},\delta^{*})$-DP for the same hyperparameter values $b$, $\sigma$ and $T$, giving some GDP parameter $\mu^{*}$. Along the GDP privacy profile determined by $\mu^{*}$, we find $(\varepsilon_{1},\delta(\varepsilon_{1}))$ that optimizes the bound of Eq. (F.1), and then evaluate the DP-SGD-$\delta$ using that same value $\varepsilon_{1}$, giving some value $\widehat{\delta}(\varepsilon_{1})$. Taking the maximum of $\delta(\varepsilon_{1})$ and $\widehat{\delta}(\varepsilon_{1})$ for the evaluation of $\widehat{\varepsilon}$ in the bound of Eq. (F.1) will then give a privacy profile that bounds the DP-guarantees of the hyperparameter tuning of both DP-SGD and NoisyCGD.
Overall, in case the batch size, number of epochs and $\sigma$ are fixed, in addition to the learning rate $\eta$, we have in all alternatives only one hyperparameter to tune: the hidden-layer width $W$ for the ReLU networks and the number of hyperplanes $P$ for the convex models.
Table 2: Model accuracies vs. $\varepsilon$-values for the DP hyperparameter tuning algorithm. The number of candidate models $K$ is Poisson distributed with mean $m=20$. The iterative methods generally have comparable accuracy and score better than SSP – similar to the conclusions from Table 1.
MNIST CIFAR-10
$\varepsilon=2.88$ $\varepsilon=8.91$ $\varepsilon=2.88$ $\varepsilon=8.91$
Sufficient Statistics Perturbation (Convex Approx.) $62.8_{{\color[rgb]{.5,.5,.5}\pm.4}}$ $71.9_{{\color[rgb]{.5,.5,.5}\pm.1}}$ $21.7_{{\color[rgb]{.5,.5,.5}\pm 1.0}}$ $24.5_{{\color[rgb]{.5,.5,.5}\pm.7}}$
Sufficient Statistics Perturbation (Random ReLU) $61.7_{{\color[rgb]{.5,.5,.5}\pm.4}}$ $68.3_{{\color[rgb]{.5,.5,.5}\pm.1}}$ $23.2_{{\color[rgb]{.5,.5,.5}\pm.5}}$ $28.5_{{\color[rgb]{.5,.5,.5}\pm.4}}$
Sufficient Statistics Perturbation (RFF) $73.5_{{\color[rgb]{.5,.5,.5}\pm.5}}$ $80.1_{{\color[rgb]{.5,.5,.5}\pm.3}}$ $25.6_{{\color[rgb]{.5,.5,.5}\pm.3}}$ $31.1_{{\color[rgb]{.5,.5,.5}\pm.1}}$
DP-SGD + Convex Approximation $92.8_{{\color[rgb]{.5,.5,.5}\pm.3}}$ $94.8_{{\color[rgb]{.5,.5,.5}\pm.2}}$ $41.6_{{\color[rgb]{.5,.5,.5}\pm.2}}$ $45.6_{{\color[rgb]{.5,.5,.5}\pm.3}}$
DP-SGD + ReLU $91.6_{{\color[rgb]{.5,.5,.5}\pm.3}}$ $94.1_{{\color[rgb]{.5,.5,.5}\pm.2}}$ $42.5_{{\color[rgb]{.5,.5,.5}\pm.2}}$ $47.2_{{\color[rgb]{.5,.5,.5}\pm.2}}$
NoisyCGD + Convex Approximation $92.4_{{\color[rgb]{.5,.5,.5}\pm.3}}$ $94.3_{{\color[rgb]{.5,.5,.5}\pm.2}}$ $41.1_{{\color[rgb]{.5,.5,.5}\pm.2}}$ $45.6_{{\color[rgb]{.5,.5,.5}\pm.3}}$
F.1 Hyperparameter Grids Used for the Experiments
@@
The hyperparameter grids for the number of random hyperplanes $P$ for the convex model and the hidden width $W$ for the ReLU network are chosen based on the GPU memory of the available machines.
@@
$\{10^{-3.0},10^{-2.5},10^{-2.0},10^{-1.5},10^{-1.0},10^{-0.5}\}.$ (F.2)
For MNIST and FashionMNIST, we tune the number of random hyperplanes $P$ over $\{32,64\}$ and for CIFAR10 over $\{16,32,64\}.$ For MNIST and FashionMNIST, we tune the hidden width $W$ of the ReLU network over $\{200,500,800\}$ and for CIFAR10 over $\{200,400\}.$
F.2 Results on DP Hyperparameter Tuning
Table 2 shows the accuracies of the best models obtained using the DP hyperparameter tuning algorithm. With the noise scale values $\sigma=5.0$ and $\sigma=15.0$, the DP-SGD trained models are $(1.33,10^{-5})$-DP and $(4.76,10^{-5})$-DP, respectively. To have similar privacy guarantees for the base models trained using NoisyCGD, we adjust the regularization constant $\lambda$ accordingly (as depicted in Section F), which leads to equal final $(\varepsilon,\delta)$-DP guarantees for the hyperparameter tuning algorithms. We see from Tables 2 that the convex models are on par in accuracy with the ReLU network. Notice also from the results of Section 5 that the learning rate tuned logistic regression cannot reach similar accuracies as the NoisyCGD trained convex model.
The results of Table 2 were the most computationally intensive part of our experiments and were run on eight NVIDIA GeForce RTX 3090 GPUs for approximately 3 days each.
F.3 Batch Size Ablation
To justify the choice of batch size 1000 for all experiments of Section 5, we consider an ablation experiment for MNIST where we study the effect of the batch size. Figures 6 and 7 show the effect of the batch size on the final test accuracy for three alternatives, when we train all models for 400 epochs. Each point is an average of three runs, and the learning rate for each model is optimized using the grid (F.2).
Figure 6: Ablation of batch size for MNIST: test accuracy vs. batch size with $\delta=10^{-5}$ and 400 training epochs. Left: DP-SGD on the convexified model. Right: NoisyCGD on the convexified model.
@@ F.3 Batch Size Ablation
Figure 7: MNIST Comparisons: Test accuracies vs. batch size for DP-SGD applied on a ReLU network with hidden layer width 500, when $\delta=10^{-5}$ and each model is trained for 400 epochs.
Appendix G Experimental Details for Baselines
G.1 Sufficient Statistics Perturbation (SSP)
We compare the proposed method with private linear regression using Sufficient Statistics Perturbation (SSP) from Amin et al., (2023). SSP is a one-step method that computes and perturbs the sufficient statistics without requiring iterations. This has the advantage that computation only scales linearly with the number of samples and cubically with the number of features, i.e., for calculating the inverse of the sample covariance matrix. The method clips input features adaptively and adds noise to the sufficient statistics as part of achieving differential privacy. Following (Wang, , 2018), we estimate the clipping norm directly from data rather than setting it ahead of time. This simplification yields optimistic results compared to pre-determined clipping – in theory, one should set the clipping norm ahead of time. However, it allows direct evaluation of the baseline experiment across datasets without introducing another hyperparameter. This does not affect our conclusions, since no result from SSP outperforms the iterative methods in our experiments.
All the datasets are multi-class classification problems, and we run multiple one-versus-all linear regressions. For example, for MNIST with 10 classes, we run the SSP method ten times – one for each class, by predicting 1 for the specific class and 0 for all other classes. The number of features is the same as the number of features in the other comparing method. This means that for MNIST and FashionMNIST, there are either 100 random hyperplanes for the ReLU methods, or 100 random Fourier features. For CIFAR10, we use only 64 random features for comparability with the iterative methods which have computational constraints. The SSP has one hyperparameter (parameter $\rho$ in Amin et al., , 2023) and we set it to the recommended default value of $0.05$.
G.2 Logistic Regression with DP-SGD
Previous work shows that DP-SGD applied to logistic regression has simnilar results as a hidden-state analysis (Bok et al., , 2024). To this end, we compare against DP-SGD applied to logistic regression in our experiments, and provide the details of the experiment in this section.
The logistic regression model is a one-layer version of the ReLU network used in our proposed method. The model is trained using DP-SGD with a Gaussian mechanism. The learning rate is tuned with the same hyperparameter tuning method as the proposed method, c.f. Section F.
The logistic regression model uses softmax classification for multi-class prediction, defined as
$f(x)_{j}=\frac{e^{w_{j}^{\top}x+b_{j}}}{\sum_{k=1}^{J}e^{w_{k}^{\top}x+b_{k}}},$ (G.1)
where $f(x)_{j}$ is the predicted probability for class $j$, $w_{j}\in\mathbb{R}^{d}$ is the weight vector for class $j$, $b_{j}\in\mathbb{R}$ is the bias term for class $j$, and $J$ is the number of classes. The model parameters are trained jointly using the cross-entropy loss
$\ell(W,b,x,y)=-\sum_{j=1}^{J}y_{j}\log(f(x)_{j}),$ (G.2)
where $y\in\{0,1\}^{J}$ is the one-hot encoded label vector, $W=[w_{1},\ldots,w_{J}]^{\top}\in\mathbb{R}^{J\times d}$ is the weight matrix containing all class weight vectors, and $b=[b_{1},\ldots,b_{J}]^{\top}\in\mathbb{R}^{J}$ contains all bias terms.
Appendix H Additional Experimental Results on FashionMNIST
Figure 8 shows the accuracies of the best models along the training iteration of 400 epochs for the FashionMNIST experiment.
For comparison, we train a linear regression model with SSP. At $\varepsilon=4.76$, this results in $67.9_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ and $73.2_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$accuracy for Random ReLU and RFF, respectively. At $\varepsilon=1.33$, the baseline accuracies are $59.9_{{\color[rgb]{.5,.5,.5}\pm 1.0}}$ and $66.4_{{\color[rgb]{.5,.5,.5}\pm 1.1}}$%, respectively.
Figure 8: FashionMNIST: Test accuracies vs. the spent privacy budget $\varepsilon$, when each model is trained for 400 epochs. The model is a 2-layer ReLU network. Similar to Figures 2 and 3, NoisyCGD and DP-SGD achieve comparable accuracies and score a lot higher than baseline regression models.
Appendix I Further Motivation for NoisyCGD Analysis: Evaluation of Shuffling Amplification Bounds
When using disjoint batches of data, a common way to obtain rigorous DP guarantees is to use shuffling amplification (Feldman et al., , 2021). However, as shown by Chua et al., 2024a ; Chua et al., 2024b , the data shuffling combined with disjoint batches leads to an inferior privacy-utility trade-off compared to random mini-batch sampling.
To illustrate the $(\varepsilon,\delta)$-DP shuffling amplification bounds of (Thm. 3.8 Feldman et al., , 2021) for DP-SGD with disjoint mini-bathces, we consider a setting in one of our experiments where we use noise parameter $\sigma=5.0$. Similarly to the experiments of Chua et al., 2024b , we use the numerical method presented in Feldman et al., (2021) to accurately compute the shuffling bounds. Further, we combine the bounds off (Thm. 3.8 Feldman et al., , 2021) with the improved bounds of (Thm. 3.1 Feldman et al., , 2023)
In our experiments of Section 5, we use batch size of 1000 which translates to 50 or 60 disjoint batches per epoch. For computing the shuffling bounds using (Thm. 3.8 Feldman et al., , 2021), one sees that 60 is too small number of batches for the conditions of the analysis to hold. We also see that the privacy guarantee clearly improves as the number of batches grows (see, e.g., the comparisons by Chua et al., 2024b, ). To obtain a lower bound for the $(\varepsilon,\delta)$-DP upper bound, we consider 1000 batches per epoch instead.
The comparison to the bounds of the Gaussian mechanism (i.e., using parallel composition) is depicted in Fig. 9 which shows that the privacy guarantees of shuffling are worse than the privacy bounds of the Gaussian mechanism. This indicates that the privacy-utility trade-offs would be inferior when using data shuffling and shuffling amplification for the DP guarantees, when compared to our approach which allows the hidden state analysis of NoisyCGD.
Figure 9: $(\varepsilon,\delta)$-DP guarantees for a single epoch of training when using 1000 disjoint batches and noise parameter $\sigma=5.0$ obtained using the shuffling amplification of (Thm 3.8, Feldman et al., , 2021). In experiments, we use 50 or 60 batches per epoch, in which case the DP guarantees are be even worse.
Appendix J Summary of Notation
  • $\ell(v,x_{j},y_{j})$: Sample-wise loss;
  • $B_{j}$: Mini-batch at iteration $j$;
  • $\beta$: smoothness parameter, Lemma 3.1 – a function $f$ is $\beta$-smooth if $\nabla f$ is $\beta$-Lipschitz;
  • $C>0$ denotes the clipping constant, $\text{clip}(\cdot,C)$ the clipping function that clips gradients to have the 2-norm at most $C$;
  • $d$: number of features of each data sample – in other words, each data is $x\in\mathbb{R}^{n\times d}$;
  • $\delta$: part of the definition of Differential Privacy – set to $10^{-5}$ in all experiments;
  • $D\sim D^{\prime}$: two adjacent datasets – datasets are adjacent when one sample is different;
  • $\varepsilon$: part of the definition of Differential Privacy – generally a number between one and ten;
  • $E$: number of (training) epochs;
  • $\eta$: learning rate;
  • $H_{\alpha}(P||Q)=\int\max\{P(t)-\alpha\cdot Q(t),0\}\,\hskip 1.0pt{\rm d}\hskip 0.5ptt.$: Hockey-stick divergence;
  • $J$: number of classes in multi-class model
  • $\kappa$: Condition number used in convex approximation analysis;
  • $f\,:\,\mathbb{R}^{d}\rightarrow\mathbb{R}$: one-layer MLP that maps $d$ dimensional input features to real valued prediction;
  • $\lambda$: parameter for strong convexity – a function $f(\cdot)$ is $\lambda$-strongly convex if the function $g(x)=f(x)-\frac{\lambda}{2}\left\lVert x\right\rVert_{2}^{2}$ is convex; note that, with overloading notation, $\lambda_{\max}$ and $\lambda_{\min}$ denote the largest and smallest eigenvalue of a matrix;
  • $\Lambda_{i}$: Diagonal boolean matrix for $i$th ReLU activation pattern;
  • $L$: gradient sensitivity, see definition 2.6;
  • $\mu$: Privacy parameter in Gaussian DP;
[…]
  • $\mathcal{L}\big{(}v,X,y\big{)}$: loss function for parameter vector $v$, matrix $X$ containing multiple examples $x_{j}$ and label vector $y$ enumerating labels $y_{j}$;
  • $\mathds{1}(\zeta)=1$, Iverson bracket, which has value 1 if $\zeta\geq 0$ and 0 otherwise – is applied elementwise in the vector context;
  • $M=\left|\mathcal{D}_{X}\right|$: the number of possible ReLU activation patterns in a two-layer ReLU network, i.e., the number of regions in a partition of $\mathbb{R}^{d}$ by hyperplanes that pass through the origin and are perpendicular to the rows of the data matrix $X\in\mathbb{R}^{n\times d}$.
  • $n$: number of training samples – introduced in Section 2;
  • $\widetilde{O}(\cdot)$ follows the well-known Big-O notation and omits logarithmic factors;
  • $P$: either the number of randomly sampled normal vectors (or hyperplanes) for the convex approximation of the ReLU network, or the number of Fourier components for the RFF model;
  • $\Pi_{\mathcal{C}}(v)=\mathrm{argmin}_{\theta\in\mathcal{C}}\left\lVert\theta-v\right\rVert_{2}$: projection of vector $v$ onto the set $\mathcal{C}$;
  • $r=\mathrm{rank}(X)$: rank of a matrix;
  • $v$: Vector of trainable parameters in convex approximation;
  • $v_{i}\in\mathbb{R}^{d}$: learnable $d$-dimensional vector. These lie in the cone $\mathcal{V}_{i}$ and are collectively named $v=\{v_{i}\}_{i=1}^{P}$;
  • $X\in\mathbb{R}^{n\times d}$: the matrix of the feature vectors, i.e., $X^{T}=\begin{bmatrix}x_{1}&\ldots&x_{n}\end{bmatrix}$
  • $y\in\mathbb{R}^{n}$: the vector of labels – depending on the model, treated as a real number or integer;
  • $Z_{j}\sim\mathcal{N}(0,\sigma^{2}I_{d})$: i.i.d. standard Gaussian noise.